OpenStudy (anonymous):

Simplify the following expression. 2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

combine the like terms. if I had 2 apples and you gave me 3 more apples I would have 5 apples. x's are apples ;)

OpenStudy (mertsj):

2x-2x=0 -2y-y=-3y 5z+3z=8z So we have: -3y+8z